OpenStudy (anonymous):

Please help! Medal will be given. -- Solve. 36 – 3x = 9 A. –12 B. –9 C. 9 D. 15

OpenStudy (paki):

you want to solve it for "X"...

OpenStudy (anonymous):

Yessir

OpenStudy (paki):

ok... leave "x" on one side and combine the like terms....

OpenStudy (anonymous):

Hmm so that'll be 36 and 9?

OpenStudy (anonymous):

Or?

OpenStudy (paki):

36 – 3x = 9 -3x=9-36 solve now....

OpenStudy (paki):

please respond here, so that i may guide you,,,

OpenStudy (anonymous):

This is so hard I really don't know how to do it :/ Is it then -3*9 ?

OpenStudy (anonymous):

Is the answer D?

OpenStudy (paki):

no, do like so... -3x=9-36 -3x=-27 or x=9

OpenStudy (paki):

not D, C...

OpenStudy (anonymous):

Hmm I can see how you got x=9 as the answer, thank you!

OpenStudy (paki):

great to hear that you got the point.... and pleasure...

OpenStudy (anonymous):

Can I come to you with anymore math questions I may have? :)

OpenStudy (paki):

yeah sure... tag me over there... i will guide you...

OpenStudy (anonymous):

Okay I will that :)

OpenStudy (paki):

ok